OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2. y − 6 =1/2 (x + 3) y =1/2 x − 6 y + 3 =1/2 (x − 6) x − 2y = 12

OpenStudy (jackellyn):

y-y1=m(x-x1) substitute (6,-3) y-(-3)=1/2(x-6) simplify y+3=1/2(x-6)